# User Simulation in the Era of Generative AI: User Modeling, Synthetic Data Generation, and System Evaluation

### v1

@@ Abstract.
User simulation is an emerging interdisciplinary topic with multiple critical applications in the era of Generative AI. It involves creating an intelligent agent that mimics the actions of a human user interacting with an AI system, enabling researchers to model and analyze user behaviour, generate synthetic data for training, and evaluate interactive AI systems in a controlled and reproducible manner. User simulation has profound implications for diverse fields and plays a vital role in the pursuit of Artificial General Intelligence. This paper provides an overview of user simulation, highlighting its key applications, connections to various disciplines, and outlining future research directions to advance this increasingly important technology.
@@ 1. Introduction
As a research topic, user simulation is inherently interdisciplinary, intersecting with diverse fields both within and beyond computer science. For example, it draws upon concepts from psychology, economics, and human-computer interaction to create accurate and representative models of user behaviour (Balog and Zhai, 2024b). The recent success of large language models (LLMs) has made it possible to simulate complex user actions and led to their widespread use in simulation tasks across different domains and application scenarios, including generating realistic conversations for dialogue systems (Kim et al., 2021), performing automatic relevance assessment of search results (Thomas et al., 2024), simulating specific human subpopulations in social science research (Argyle et al., 2023), and simulating the behaviour of communities (Park et al., 2023). Realizing the potential of user simulation, there has been a surge of interest and activity in this area, evidenced by the growing number of related workshops (Balog et al., 2022; Schaer et al., 2024) and tutorials (Balog and Zhai, 2024a; Deng et al., 2024).
{mdframed}
[roundcorner=5pt] Key highlights
  • User simulation, using intelligent agents to mimic real users, provides a controlled environment for training, testing, and refining interactive AI systems.
  • User simulation requires the creation of a mathematical model of the user, serving as a formal hypothesis about their behaviour.
  • User simulation intersects with psychology, economics, and human-computer interaction, leveraging these fields to create accurate models of user behaviour.
  • Large language models (LLMs) are increasingly used in diverse simulation tasks, from chatbot dialogues to customer interactions and community behaviour.
  • The overarching goal of developing a realistic user simulator is, in many ways, aligned with the pursuit of Artificial General Intelligence (AGI).
This article aims to synthesize research dispersed across various fields, review the current state of the art, and highlight potential future directions. Specifically, we provide a brief historical review of how simulation techniques can be employed for (1) user behaviour modeling, by creating realistic simulations of how users interact with a system when attempting to complete some task (Section 3), (2) data augmentation, by generating synthetic user interactions to improve the training of machine learning models (Section 4), and (3) system evaluation, by measuring the perceived utility and cost/effort from a user’s perspective when completing a task (Section 5). Next, we broaden our perspective on user simulation to explore its role in integrating multiple research fields (Section 6). Finally, we discuss the role user simulation may play in the quest for Artificial General Intelligence (AGI), the ultimate goal of developing AI systems with human-like intelligence (Section 7).
@@ 2. User Simulation
What exactly do we mean by user simulation? In essence, it involves creating an intelligent agent that mimics how a user interacts with a system. This agent can be built based on models/algorithms/rules and any knowledge we have about the user (their behaviour, knowledge, etc.). Crucially, the agent can be parameterized to simulate a diverse range of users with varying characteristics.
@@
Task ($T$) System ($S$) User information ($U$) Actions ($\mathcal{A}$)
@@ 2.1. Definition
User simulation is the process of modeling a user’s behaviour and decision-making patterns within an interactive system, specifically designed to mimic and predict how a user will act in various interaction contexts or scenarios related to completing a task. To effectively simulate a user’s behaviour within an interactive system, configuration variables that influence this behaviour must be defined:
  • Task ($T$): A user’s behaviour varies according to nature of the user’s task. Tasks vary in complexity, and different tasks require different types and levels of interaction, decision-making processes, and completion strategies.
  • System ($S$): A user’s behaviour depends on the system they interact with. This includes the system’s functionality, user interface, and overall usability and support for task goals. It is the system that dictates the types of possible actions, denoted as $\mathcal{A}$, that a user can perform at any given point during their interactions.
  • User information ($U$): Different users may behave differently when completing the same task using the same system. Simulations must account for variations in individual user characteristics such as age, technical proficiency, preferences, and cognitive styles.
With these variables defined, the task of user simulation can be stated as the following computational problem:
Given the variables $T$, $S$, and $U$, the goal of user simulation is to create an agent that can simulate every action that user $U$ may take when attempting to complete task $T$ using system $S$.
This problem involves developing an intelligent agent (i.e., a user simulation agent) that can dynamically generate user actions, reflecting the behavioural patterns and decision-making processes of a user, based on a specific task context. The main technical challenge is to compute a policy function: $\pi:\mathcal{S}\rightarrow A$, where $\mathcal{S}=(T,U,S,H)$ represents the current state, encompassing information about the task $T$, system $S$, user $U$, as well as the history of previous interactions $H$ (including the actions taken by the user, the responses provided by the system, and any other relevant events that have occurred during the interaction), and $A\in\mathcal{A}$ is the action taken by the (simulated) user. Once an action is taken, the current state would be “transitioned” to a new state, leading to a new decision problem as to which action to be taken based on the new state. The choice of computational model for the policy (e.g., rule-based, probabilistic, or machine-learned algorithm) is influenced by the nature of the task, system, and user information, leading to many variations of approaches to user simulation. However, in general, such a user simulation agent can be modeled based on the Markov Decison Process (MDP) with actions defined on each potential state (see (Balog and Zhai, 2024b) for more discussion of the framework and how existing approaches can be framed as instations of the MDP framework).
@@ 2.2. Scope
User simulation encompasses a wide spectrum, ranging from predicting single actions to modeling complex behaviour across multiple tasks. In our formulation, this scope is primarily determined by how the task information ($T$) is defined. At one end of the spectrum, $T$ might represent a very specific interaction context, such as predicting whether a user would click on a particular search result snippet. Here, the focus is on simulating a single, isolated action. Moving along the spectrum, $T$ could encompass a sequence of actions within a given context, such as reformulating search queries within a search session, requiring the model to consider dependencies between actions. Further expanding the scope, $T$ might represent an entire task, such as finding information on a particular topic or completing a purchase, where the simulation would involve multiple sequences of interactions. Finally, at the broadest level, $T$ could encompass a user’s general preferences and behaviour across various tasks, necessitating models that capture long-term patterns and adapt to different contexts. Thus, by varying the granularity and breadth of $T$, our formulation allows for user simulations in a wide range of application scenarios at different levels of complexity. Table 1 lists specific examples of user simulation for tasks.
2.3. Approaches
The problem of simulating the action a user takes in a given context (represented by the policy function $\pi$) can often be framed as a classification problem when there is a relatively small set of actions to choose from; for example, simulation of a user’s clicking action may be framed as a binary classification problem, where the algorithm would predict whether the simulated user would click on a specific item. When there are potentially infinitely many actions to choose from (e.g., when formulating a query, any valid query would be potentially an option), simulating user actions becomes significantly more challenging. The vast action space can lead to computational complexity and makes it difficult to define a meaningful probability distribution over possible actions. In practice, we often make assumptions to restrict the number of actions to be considered when simulating those actions (e.g., assuming the user will only use keywords from a predefined vocabulary, or that the query will follow a specific structure). However, other approaches, such as using generative models, may also be employed to handle large or infinite action spaces more effectively.
With the problem framed as a classification problem, different approaches generally vary in how they perform the classification (equivalently prediction) task. At a high level, we can distinguish two broad approaches: model-based and data-driven.
@@ 2.5. Requirements and Desiderata
Effective user simulators depend on a range of carefully considered properties, including:
@@
  • Interpretability: While not strictly a requirement, interpretability is a highly desirable property. Interpretability means that the simulated behaviour can be understood and adjusted through controllable parameters. This allows researchers to (1) understand why the simulator produced certain behaviours and (2) investigate how changes in specific parameters influence the behaviour of users. Since user behaviour and preferences vary significantly across users, interpretability is crucial for understanding which real-world users can be expected to produce results similar to the simulations.
@@
The good news is that simulation does not need to be perfect in order to be useful in a specific application context. Note that evaluation methodologies, even in well-established areas like search engine ranking, are inherently imperfect. For example, relevance judgments in almost all test collections are incomplete. However, research suggests that the relative performance of different retrieval systems tend not be affected much by this incompleteness (Voorhees, 2000). In fact, creating a “perfect” user simulator, i.e., one that flawlessly replicates human behaviour across all possible tasks and contexts, is likely an AI-complete problem, on par with achieving Artificial General Intelligence (AGI, cf. Section 7).
@@ 3. User Simulation for User Modeling
Lastly, as a valuable tool for user studies, a formal user simulator can facilitate various comparative analyses in association with other relevant user variables. For instance, by partitioning a large user population along various dimensions (e.g., age, occupation, or geo-locations), user simulation models can be employed to conduct comparative analyses among different subgroups to understand their commonalities and differences as reflected in the observed user data such as interaction logs. Similarly, any variable related to an AI system can also be varied to examine how a user’s behaviour may vary depending on system-related variables; for example, the device and time available can profoundly affect what kind of interactive searching behaviour can be successful (Baskaya et al., 2012).
With a simulator of an individual user as a building block, it is also possible to simulate a community of people, which, theoretically speaking, can be any set of individuals that are connected in some way. By representing the behaviour of a community as an aggregate function of the simulated behaviours of its individual users, we can study its overall dynamics and evolution over time. For example, a community may include the entire user base of an application such as a recommender system, and user simulation can be used to study the long-term impact of, for instance, how content creators’ strategic behaviour in competing for user engagement (Yao et al., 2023) or the tendency of users choosing more popular items (Hazrati and Ricci, 2024) influences the overall content ecosystem and user experience within the platform. Modeling a community introduces the complex challenge of simulating how users would interact with each other—an area LLMs will likely play an important role (Park et al., 2023).
@@ 6. User Simulation as an Interdisciplinary Research Field
Intelligent Agents
A sophisticated user simulator can be regarded as an intelligent agent interacting within a system’s environment. Consequently, techniques used for developing intelligent agents can also prove valuable in the construction of user simulators. More specifically, multi-agent systems are particularly well-suited for simulating communities of users, as they enable the modeling of complex interactions among multiple individuals within the simulated environment (Uhrmacher and Weyns, 2009). A perfect user simulation agent can also be regarded as a basis to implement a human digital twin, which has widespread applications (Wang et al., 2024b). Conversely, building upon user simulation research, intelligent agents can benefit from even more precise modeling of individual behaviour.
Machine Learning
@@
HCI and Psychology
Psychology and HCI research are fundamental for building realistic user simulators. Specifically, this can include insights gained from user studies, as well as well-motivated and explicit computational cognitive models that explain certain behaviours, such as navigation behaviour on the Web (Fu and Pirolli, 2007). At the same time, an interpretable user simulator can be instrumental in analyzing user interaction data for deeper behavioural understanding. Existing research has employed various models to analyze interaction logs; for example, mining search engine logs reveals benefits of users following search trails versus jumping directly to destination pages (White and Huang, 2010). Furthermore, simulators can serve as testable “computational models” of users, enabling the formulation and validation of hypotheses about user behaviour.
Software Systems

### v2

@@ Abstract.
User simulation is an emerging interdisciplinary topic with multiple critical applications in the era of Generative AI. It involves creating an intelligent agent that mimics the actions of a human user interacting with an AI system, enabling researchers to model and analyze user behaviour, generate synthetic data for training, and evaluate interactive AI systems in a controlled and reproducible manner. Because of its broad scope, research on this topic currently remains scattered across artificial intelligence, human-computer interaction, information science, computational social science, and psychology. To address this fragmented landscape of current research, this article presents a foundational synthesis. We highlight the paradigm shift from traditional predictive models to modern generative approaches, and explicitly frame critical ethical considerations—demonstrating how controlled simulation serves not merely as a risk vector for bias, but as a powerful, proactive tool to ensure fair representation and system safety. Furthermore, we establish the theoretical connection between user simulation and the pursuit of Artificial General Intelligence, arguing that realistic simulators are indispensable catalysts for overcoming critical data and evaluation bottlenecks and optimizing personalization. Ultimately, we propose a practical, self-sustaining innovation ecosystem bridging academia and industry to advance this increasingly important technology.
@@ 1. Introduction
As a research topic, user simulation is inherently interdisciplinary, intersecting with diverse fields both within and beyond computer science. For example, it draws upon concepts from psychology, economics, and human-computer interaction to create accurate and representative models of user behaviour (Balog and Zhai, 2024b). The recent success of large language models (LLMs) has made it possible to simulate complex user actions and led to their widespread use in simulation tasks across different domains and application scenarios, including generating realistic conversations for dialogue systems (Kim et al., 2021), performing automatic relevance assessment of search results (Thomas et al., 2024), simulating specific human subpopulations in social science research (Argyle et al., 2023), and simulating the behaviour of communities (Park et al., 2023). Realizing the potential of user simulation, there has been a surge of interest and activity in this area, evidenced by the growing number of related workshops (Balog et al., 2022; Breuer et al., 2024; Schaer et al., 2025) and tutorials (Balog and Zhai, 2024a; Deng et al., 2024; Balog et al., 2025; Zerhoudi et al., 2026).
This article aims to synthesize research dispersed across various fields, review the current state of the art, and highlight potential future directions. Rather than proposing new models or architectures, our primary contribution is to provide a foundational overview. By connecting fragmented efforts across disciplines, we aim to establish a shared conceptual framework. Specifically, after outlining the paradigm shift from traditional predictive modeling to modern generative approaches, we provide a brief historical review of how simulation techniques can be employed for (1) user behaviour modeling, by creating realistic simulations of how users interact with a system when attempting to complete some task, (2) data augmentation, by generating synthetic user interactions to improve the training of machine learning models, and (3) system evaluation, by measuring the perceived utility and cost/effort from a user’s perspective when completing a task. Next, we broaden our perspective on user simulation to explore its role in integrating multiple research fields. Following this, we address the critical ethical considerations inherent in user simulation. Crucially, we reframe these challenges to demonstrate how simulation serves not merely as a risk vector for bias, but as a powerful, proactive tool to debias training pipelines, ensure fair subpopulation representation, and systematically stress-test safety guardrails. Finally, we discuss the role user simulation may play in the quest for Artificial General Intelligence (AGI), the ultimate goal of developing AI systems with human-like intelligence. To ensure the continuous and scalable advancement of this technology, we conclude by proposing a practical, self-sustaining innovation ecosystem that bridges academia and industry.
@@ 2. User Simulation
What exactly do we mean by user simulation? In essence, it involves creating an intelligent agent that mimics how a user interacts with a system. This agent can be built based on models/algorithms/rules and any knowledge we have about the user (their behaviour, knowledge, etc.). Crucially, the agent can be parameterized to simulate a diverse range of users with varying characteristics. While historically constrained by narrow, task-specific action spaces, the advent of Generative AI has fundamentally transformed how these agents are constructed, a paradigm shift toward more holistic user simulation we explore in detail below.
@@
Task context System environment User characteristics Action space
@@ 2.1. Definition
User simulation is the process of modeling a user’s behaviour and decision-making patterns within an interactive system, specifically designed to mimic and predict how a user will act in various interaction contexts or scenarios related to completing a task. To effectively simulate a user’s behaviour within an interactive system, three core dimensions that influence this behaviour must be defined:
  • Task context: A user’s behaviour varies according to nature of the user’s task. Tasks vary in complexity, and different tasks require different types and levels of interaction, decision-making processes, and completion strategies.
  • System environment: A user’s behaviour depends on the system they interact with. This includes the system’s functionality, user interface, and overall usability and support for task goals. It is the system that dictates the types of possible actions that a user can perform at any given point during their interactions.
  • User characteristics: Different users may behave differently when completing the same task using the same system. Simulations must account for variations in individual user characteristics such as age, technical proficiency, preferences, and cognitive styles.
With these dimensions defined, the task of user simulation can be stated as the following computational problem:
The goal of user simulation is to create an agent that can simulate every action a specific user may take when attempting to complete a specific task using a specific system.
This problem involves developing an intelligent agent (i.e., a user simulation agent) that can dynamically generate user actions, reflecting the behavioural patterns and decision-making processes of a user, based on a specific system environment and task context. The main technical challenge is to compute a decision-making policy. This policy must evaluate the current state—encompassing information about the task, the user’s traits, the system environment, and the history of previous interactions—to determine the next most plausible action. Once an action is taken, the environment transitions to a new state, leading to a new decision problem. While the choice of computational model for the policy (e.g., rule-based, probabilistic, or machine-learned algorithm) varies, such user simulation agents can generally be modeled based on the Markov Decison Process (MDP) framework (see (Balog and Zhai, 2024b) for a deeper discussion on instantiating existing approaches within the MDP framework).
@@ 2.2. Scope
User simulation encompasses a wide spectrum, ranging from predicting single actions to modeling complex behaviour across multiple tasks. In our formulation, this scope is primarily determined by how the task context is defined. At one end of the spectrum, this context might represent a very specific interaction, such as predicting whether a user would click on a particular search result snippet. Here, the focus is on simulating a single, isolated action. Moving along the spectrum, the context could encompass a sequence of actions within a given session, such as reformulating search queries, requiring the model to consider dependencies between actions. Further expanding the scope, it might represent an entire overarching task, such as finding information on a particular topic or completing a purchase, where the simulation would involve multiple sequences of interactions. Finally, at the broadest level, the scope could encompass a user’s general preferences and behaviour across various tasks, necessitating models that capture long-term patterns and adapt to different environments. Thus, by varying the granularity and breadth of the task context, our formulation allows for user simulations in a wide range of application scenarios at different levels of complexity. Table 1 lists specific examples of user simulation for tasks.
2.3. Traditional Approaches
The problem of simulating the action a user takes in a given context (represented by the decision policy) can often be framed as a classification problem when there is a relatively small set of actions to choose from; for example, simulation of a user’s clicking action may be framed as a binary classification problem, where the algorithm would predict whether the simulated user would click on a specific item. With the problem framed as a classification problem, different approaches generally vary in how they perform the classification (equivalently prediction) task. At a high level, we can distinguish two broad approaches: model-based and data-driven.
@@ 2.3. Traditional Approaches
Table 2. Comparison of traditional user simulation approaches versus the Generative AI paradigm.
Model-based Data-driven Generative AI
Core mechanism Explicit rules or interpretable probabilistic models Supervised machine learning (e.g., deep neural networks) Foundation models (LLMs) pre-trained on broad world knowledge
Action space Highly constrained (predictive selection from predefined choices) Constrained to specific patterns present in the training data Open-ended (natural language, code generation, complex sequential actions)
Primary strengths High interpretability; allows precise counterfactual tuning and parameter variation High predictive validity; effectively captures complex implicit patterns directly from interaction logs Zero-shot generalization; capable of complex, multi-turn interactions; easily parameterizable via prompting
Limitations & weaknesses Rigid; heavily relies on manual feature engineering; struggles to capture complex, nuanced human behaviour “Black box” nature with low interpretability; requires large amounts of interaction data; lacks zero-shot generalization Prone to behavioural hallucination; difficult to strictly control cognitive constraints (e.g., bounded rationality)
When there are potentially infinitely many actions to choose from (e.g., when formulating a query, any valid query would be potentially an option), simulating user actions becomes significantly more challenging for traditional approaches. The vast action space can lead to computational complexity and makes it difficult to define a meaningful probability distribution over possible actions. In practice, we often make assumptions to restrict the number of actions to be considered when simulating those actions (e.g., assuming the user will only use keywords from a predefined vocabulary, or that the query will follow a specific structure). However, modern approaches, notably the large generative models driving the current wave of Generative AI, can handle large or infinite action spaces much more effectively.
2.4. The Generative AI Paradigm Shift
Building on this capability to handle open-ended action spaces, the advent of Generative AI introduces a broader technical paradigm shift: the transition from predictive to generative simulation. Historically, user simulation predominantly relied on predictive classification—such as predicting whether a user will click a specific link or assign a specific rating. These traditional models require task-specific feature engineering, struggle to generalize beyond their training data, and are entirely incapable of handling open-ended interactions. In contrast, rather than learning isolated behaviours from scratch, LLMs bring foundational world knowledge, semantic understanding, and zero-shot generalization to the simulation environment. This allows GenAI-based simulators to operate within infinite, unconstrained action spaces—such as formulating natural language queries, engaging in complex multi-turn dialogues, or generating novel code—which were previously impossible to simulate at scale. Consequently, GenAI elevates user simulation from a narrow, task-specific evaluation tool into a versatile, foundational building block capable of modeling complex cognitive tasks. To summarize this evolution, Table 2 provides a comprehensive comparison of traditional model-based and data-driven approaches against the emerging Generative AI paradigm, highlighting their respective core mechanisms, action spaces, strengths, and limitations.
@@ 2.6. Requirements and Desiderata
Table 3. Desirable properties of user simulators and their inherent trade-offs and challenges.
Property Definition Key trade-offs & challenges
Validity Simulated behaviours align with empirical observations of real user behaviour in similar contexts High predictive validity often comes at the cost of interpretability, as highly accurate data-driven models are frequently opaque
Interpretability The simulated behaviour can be understood and adjusted through explicitly controllable parameters Often limits the complexity and predictive power of the model; highly interpretable models may fail to capture nuanced user interactions
Cognitive plausibility Decision-making processes are grounded in theories or models of human cognition, preventing arbitrary actions Translating complex psychological theories into computable algorithms is fundamentally difficult; current LLMs mimic idealized reasoning rather than actual human cognitive constraints (e.g., bounded rationality, fatigue, and limited memory)
Variation Incorporates natural human variability and occasional outliers rather than merely replicating average behaviour Difficult to calibrate effectively; excessive variation can easily lead to arbitrary, unrealistic, or unsafe simulated behaviours
Adaptability User simulator agents can learn from their interactions, updating their expectations and adjusting behaviour accordingly over time Requires complex, dynamic memory mechanisms (e.g., maintaining evolving knowledge states) and significantly increases overall simulation complexity
Designing an effective user simulator requires balancing several complex, and often competing, requirements. Table 3 summarizes the core desirable properties of user simulators, outlining their definitions alongside the inherent trade-offs and challenges involved in achieving them.
@@
  • Interpretability: While not strictly a requirement, interpretability is a highly desirable property. Interpretability means that the simulated behaviour can be understood and adjusted through controllable parameters. This allows researchers to (1) understand why the simulator produced certain behaviours, (2) investigate how changes in specific parameters influence the behaviour of users, and (3) systematically introduce controlled variation into the simulated population. Since user behaviour and preferences vary significantly across users, interpretability is crucial for understanding which real-world users can be expected to produce results similar to the simulations.
@@
The good news is that simulation does not need to be perfect in order to be useful in a specific application context. Note that evaluation methodologies, even in well-established areas like search engine ranking, are inherently imperfect. For example, relevance judgments in almost all test collections are incomplete. However, research suggests that the relative performance of different retrieval systems tend not be affected much by this incompleteness (Voorhees, 2000). In fact, creating a “perfect” user simulator, i.e., one that flawlessly replicates human behaviour across all possible tasks and contexts, is likely an AI-complete problem, on par with achieving Artificial General Intelligence (AGI).
@@ 3. User Simulation for User Modeling
Lastly, as a valuable tool for user studies, a formal user simulator can facilitate various comparative analyses in association with any relevant user variables. For instance, by partitioning a large user population along various dimensions (e.g., age, occupation, or geo-locations), user simulation models can be employed to conduct comparative analyses among different subgroups to understand their commonalities and differences as reflected in the observed user data such as interaction logs. Similarly, any variable related to an AI system can also be varied to examine how a user’s behaviour may vary depending on system-related variables; for example, the device and time available can profoundly affect what kind of interactive searching behaviour can be successful (Baskaya et al., 2012).
With a simulator of an individual user as a building block, it is also possible to simulate a community of people, which, theoretically speaking, can be any set of individuals that are connected in some way. By representing the behaviour of a community as an aggregate function of the simulated behaviours of its individual users, we can study its overall dynamics and evolution over time. For example, a community may include the entire user base of an application such as a recommender system, and user simulation can be used to study the long-term impact of, for instance, how content creators’ strategic behaviour in competing for user engagement (Yao et al., 2023a) or the tendency of users choosing more popular items (Hazrati and Ricci, 2024) influences the overall content ecosystem and user experience within the platform. Modeling a community introduces the complex challenge of simulating how users would interact with each other—an area LLMs will likely play an important role (Park et al., 2023).
@@ 6. User Simulation as an Interdisciplinary Research Field
6.1. Intelligent Agents
A sophisticated user simulator can be regarded as an intelligent agent interacting within a system’s environment. Consequently, techniques used for developing intelligent agents can also prove valuable in the construction of user simulators. More specifically, multi-agent systems are particularly well-suited for simulating communities of users, as they enable the modeling of complex interactions among multiple individuals within the simulated environment (Uhrmacher and Weyns, 2009). A perfect user simulation agent can also be regarded as a basis to implement a human digital twin, which has widespread applications (Wang et al., 2024a). Conversely, building upon user simulation research, intelligent agents can benefit from even more precise modeling of individual behaviour.
6.2. Machine Learning
@@ 6.3. Knowledge Representation
6.4. HCI and Psychology
Psychology and HCI research are fundamental for building theoretically sound and realistic user simulators. Specifically, this can include insights gained from user studies, as well as well-motivated and explicit computational cognitive models that explain certain behaviours, such as navigation behaviour on the Web (Fu and Pirolli, 2007), which can directly inform the design of interpretable user simulators. At the same time, an interpretable user simulator can be instrumental in analyzing user interaction data for deeper behavioural understanding. Existing research has employed various models to analyze interaction logs; for example, mining search engine logs reveals benefits of users following search trails versus jumping directly to destination pages (White and Huang, 2010). Furthermore, simulators can serve as testable “computational models” of users, enabling the formulation and validation of hypotheses about user behaviour.
6.5. Software Systems